Paragraph Read each question carefully and provide the requested information in the provided space, making sure you show any formulas, calculations, graphs, and answers (label everything correctly). Styles The market for cigars is characterized by Qd-10-0.25P and QS-0.15P, where P İsthe price per box of cigars and Q measures boxes per hour. a. What is the equation for the inverse demand curve? b. What is the equation for the inverse supply curve? c. What is the choke price for the demand curve? d. What is the choke price for the supply curve? e. What is the consumer surplus before any cigar tax? f. What is the producer surplus before any cigar tax? 8 g. Draw and label the demand curve, the supply curve, the pre-tax consumer surplus and pre-tax producer surplus.
Answer #1

1.

A.

P = (1/.25)*(10-Qd)

P = 40 - 4Qd (Inverse demand equation)

B.

P = (1/.15)*Qs

C.

Choke price = price at that demand is zero

P = 40-4*0

P = $40

D.

P = (1/.15)*0

P = 0

E.

40-4Q = (1/.15)*Q

6 - .6Q = Q

Q = 6/1.6 = 3.75

P = 3.75/.15 = $25

Consumer surplus = .5*(40-25)*3.75 = $28.125

F.

Producer surplus = .5*(25-0)*3.75 = $46.875

Pl. repost other unanswered questions for their proper answers!